Project

General

Profile

Actions

20160927 mipo-experimental

Successes

  • More-or-less stable running of magnets
  • Magnet sign off - we should do this again
  • Magnet on-call - we should keep up the good work here
  • We got data!

Data taken

Only field-on data is listed. We have not got much useful field-off data due to issues listed below - Sunday and Monday were wiped out by DAQ and controls issues, until ISIS went down at 17:00 on Monday (just as things were sort of working).

140 MeV/c

Run # Time Number TOF1 triggers Number TOF2 triggers
8265 255 minutes 88k 53k
8276 80 minutes 22k
8282 30 minutes 6k

170 MeV/c

Run # Time Number TOF1 triggers Number TOF2 triggers
8277 105 minutes 23k
8280 210 minutes 145k 70k

200 MeV/c

Run # Time Number TOF1 triggers Number TOF2 triggers
8281 240 minutes 212k 70k

Issues that affected data taking

Some may be fixed

  1. Water flow is blocking now
  2. Diffuser may become blocking next week
  3. Run control instability
  4. Tracker DAQ instability
  5. DAQ central instability (is this triggered by tracker DAQ?)
  6. Alarm handler issues
  7. General controls issues
    • Trim coil current instability
    • Beam stop registered closed when not
    • Temperature readout on BL water reading incorrect value
  8. TOF dead channels
  9. Quadrature error in target - possibly user error

Alarms Sounding

  1. Diffuser: known broken equipment
  2. Heartbeat: false alarm
  3. CDB communications
  4. tracker cryo PSU voltage: known broken equipment
  5. tracker AFE temp: known broken equipment
  6. proton absorber
  7. FC insulating vacuum
  8. T2 current: false alarm
  9. M2 current: false alarm
  10. FC lHe level: known issue, chronically flaky equipment
  11. FC load cell

Updated by Rogers, Chris over 6 years ago ยท 10 revisions locked